# The Meta-Representation Hypothesis

### v3

@@ Abstract
Humans rely on high-level understandings of things, i.e., meta-representations, to engage in abstract reasoning. In complex cognitive tasks, these meta-representations help individuals abstract general rules from experience. However, constructing such meta-representations from high-dimensional observations remains a longstanding challenge for reinforcement learning (RL) agents. For instance, a well-trained agent often fails to generalize to even minor variations of the same task, such as changes in background color, while humans can easily handle. In this paper, we theoretically investigate how meta-representations contribute to the generalization ability of RL agents, demonstrating that learning meta-representations from high-dimensional observations enhance an agent’s ability to generalize across varied environments. We further hypothesize that deep mutual learning (DML) among agents can help them learn the meta-representations that capture the underlying essence of the task. Empirical results provide strong support for both our theory and hypothesis. Overall, this work provides a new perspective on the generalization of deep reinforcement learning.
Project website: Click here to view our website.
Machine Learning, ICML
@@ 1 Introduction
A meta-representation refers to a higher-order form of representation—essentially, a representation of a representation (Wilson, 2012; Redshaw, 2014). In other words, it is an abstraction that captures not just the content of an experience or concept, but how that content is represented. To illustrate, consider the saying, “There are a thousand Hamlets in a thousand people’s eyes.” Here, the text of Hamlet serves as a direct representation, whereas each reader’s interpretation of the play is a form of meta-representation—an abstract, high-level understanding of the content.
Figure 1: Pablo Picasso’s The Bull (Scott, 2019). By focusing on and exaggerating specific details, rather than trying to capture every detail realistically, artists can convey the core meaning or essence of the subject—meta-representation, more powerfully.
Humans process and integrate vast amounts of information from the real world through meta-representations, which are underlying structured information beyond the direct sensory representations of things we perceive (Figure 1). These meta-representations enable us to generalize across tasks with similar underlying semantics. For instance, once we have learned how to play a video game, we can apply the same skills even if the game’s visual presentation changes. This suggests that the ability to perform tasks is not tied to the specific visual details of the game, but to the underlying cognitive processes that abstract away from these changes. The development of abstract thinking is linked to the human prefrontal cortex (Bengtsson et al., 2009; Dumontheil, 2014), and certain inhibitory neurons further enhance the brain’s processing efficiency (Pi et al., 2013).
While humans can generalize across tasks by relying on abstract meta-representations, visual reinforcement learning (VRL) faces a significant challenge in this regard. Although well-trained agents can solve complex tasks, they often struggle to transfer their experience to new environments. Even subtle changes, such as variations in scene colors, can hinder their ability to generalize, demonstrating that their learning is overly dependent on specific visual inputs (Cobbe et al., 2019, 2020).
What makes it difficult for reinforcement learning agents to generalize? How can these agents develop the ability to construct meta-representations just like humans?
Our central theory, presented in Algorithm 1, assumes the existence of several Markov Decision Processes (MDPs) sharing an underlying MDP. Imagine the scenario of building a reinforcement learning benchmark to test the generalization performance of algorithms. We would first implement the core code for the underlying MDP $\mathcal{M}$, which reflects the intrinsic properties of the task. Then we randomly initialize a rendering function $f$, which obfuscates the underlying state $s_{t}$ into the agent’s observation $o_{t}=f(s_{t})$, akin to how different schools of painters might depict the same scene in various styles. To achieve good generalization, the agent must learn to ignore the interference from $f$. In this scenario, learning the meta-representation means that the agent has learned to perceive beyond the noisy observation $o_{t}$ and grasp the true underlying state $s_{t}$. This process is far more challenging than simply achieving high performance during training, as it requires the agent to filter out the noise and focus on the core task structure.
Algorithm 1 MDP Generator
1:  Initialize: Underlying MDP $\mathcal{M}$ and behavior policy $\pi$
2:  while collecting data do
3:     Randomly initialize a rendering function $f$
4:     Underlying initial state $s_{0}\sim\mathcal{M}$
5:     for $t=0$ to $T$ do
6:        The noisy observation $o_{t}=f(s_{t})$
7:        Choose action $a_{t}\sim\pi(\cdot|o_{t})$
8:        Update environment $r_{t},s_{t+1}\sim\mathcal{M}(s_{t},a_{t})$
9:        Store data $(o_{t},a_{t},r_{t})$
10:     end for
11:  end while
This paper aims to develop a theory of generalization in reinforcement learning, with a particular focus on learning meta-representations that capture the essential structure of tasks beyond superficial observations. Unlike traditional approaches such as the Partially Observable Markov Decision Process (POMDP) (Murphy, 2000), which focuses primarily on the challenge of partial observability of the true state, our framework emphasizes the ability of agents to learn abstract, high-level representations beyond the noisy observations. Furthermore, our meta-representation hypothesis posits that deep mutual learning (DML) (Zhang et al., 2018b) between agents can facilitate the learning of these meta-representations, thereby improving generalization performance across different environments sharing the same underlying semantics. Extensive experiments on Procgen (Cobbe et al., 2019, 2020) support our theory and hypothesis, demonstrating that PPO (Schulman et al., 2017) with our DML-based framework achieves significant improvements over the standard PPO.
Overall, this study presents a novel perspective on improving the generalization capabilities of deep reinforcement learning, providing insights that could contribute to more adaptable and robust decision-making in diverse and dynamic environments. The main contributions of this paper are summarized as follows:
@@
  • We theoretically prove that improving the policy robustness to irrelevant features enhances generalization performance. To the best of our knowledge, we are the first to provide a rigorous proof of this intuition.
@@
  • We propose a hypothesis that deep mutual learning (DML) can facilitate the learning of meta-representations by agents, and we also provide intuitive insights to support this hypothesis.
@@ 2.1 Markov Decision Process and Generalization
$\begin{split}Q^{\mu}(s_{t},a_{t})&=\mathbb{E}_{\mu}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k})\right],\\ V^{\mu}(s_{t})&=\mathbb{E}_{a_{t}\sim\mu(\cdot|s_{t})}\left[Q^{\mu}(s_{t},a_{t})\right].\\ \end{split}$ (1)
@@
Explanation. Assumption 2.2 states that all $\mathcal{M}_{f}$ share a common underlying MDP $\mathcal{M}$, which is a formal statement of Algorithm 1.
@@
where $o_{0}^{f}\sim\rho_{f}(\cdot)$, $a_{t}\sim\pi(\cdot|o_{t}^{f})$, $r_{t}^{f}=r_{f}(o_{t}^{f},a_{t})$ and $o_{t+1}\sim\mathcal{P}_{f}(\cdot|o_{t}^{f},a_{t})$, we simplify the notation to $\tau_{f}\sim\pi$.
However, during training, the agent is only allowed to access a subset of all MDPs, which is $\left\{\mathcal{M}_{f}|f\in\mathcal{F}_{\mathrm{train}}\subset\mathcal{F}\right\}$, and then tests its generalization performance across all MDPs. Thus, denote $p_{\mathrm{train}}:\mathcal{F}_{\mathrm{train}}\mapsto[0,1]$ as the distribution of $\mathcal{F}_{\mathrm{train}}$, the agent’s training performance $\eta(\pi)$ and generalization performance $\zeta(\pi)$ can be expressed as
$\begin{split}\eta(\pi)&=\mathbb{E}_{f\sim p_{\mathrm{train}}(\cdot),\tau_{f}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{f}(o_{t}^{f},a_{t})\right],\\ \zeta(\pi)&=\mathbb{E}_{f\sim p(\cdot),\tau_{f}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{f}(o_{t}^{f},a_{t})\right].\\ \end{split}$ (4)
@@ 2.2 Deep Mutual Learning
Deep mutual learning (DML) (Zhang et al., 2018b) is a mutual distillation technique in supervised learning. Unlike the traditional teacher-student distillation strategy, DML aligns the probability distributions of multiple student networks by minimizing the KL divergence loss during training, allowing them to learn from each other. Specifically,
@@
where $\mathcal{L}_{\mathrm{SL}}$ and $\mathcal{L}_{\mathrm{KL}}$ represent the supervised learning loss and the KL divergence loss, respectively, $\alpha$ is the weight.
Using DML, the student cohort effectively pools their collective estimate of the next most likely classes. Finding out and matching the other most likely classes for each training instance according to their peers increases each student’s posterior entropy, which helps them converge to a more robust representation, leading to better generalization.
@@ 3 Theoretical Results
$\begin{split}p_{\mathrm{train}}(f)&=\frac{p(f)\cdot\mathbb{I}(f\in\mathcal{F}_{\mathrm{train}})}{Z},\\ p_{\mathrm{eval}}(f)&=\frac{p(f)\cdot\mathbb{I}(f\in\mathcal{F}_{\mathrm{eval}})}{1-Z},\\ \end{split}$ (6)
where $Z=\int_{\mathcal{F}_{\mathrm{train}}}p(f)\mathrm{d}f$ and $1-Z$ is the partition function, $\mathcal{F}_{\mathrm{eval}}=\mathcal{F}-\mathcal{F}_{\mathrm{train}}$, $\mathbb{I}(\cdot)$ denotes the indicator function.
@@
$\begin{split}\eta(\pi)&=\frac{1}{1-\gamma}\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\mu_{f}}(\cdot)\\ a\sim\mu_{f}(\cdot|s)\end{subarray}}\left[r(s,a)\right],\\ \zeta(\pi)&=\frac{1}{1-\gamma}\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p(\cdot)\\ s\sim d^{\mu_{f}}(\cdot)\\ a\sim\mu_{f}(\cdot|s)\end{subarray}}\left[r(s,a)\right].\\ \end{split}$ (7)
We can thus analyze the generalization problem using the underlying policy $\mu_{f}$. Then, define $L_{\pi}$ as the first-order approximation of $\eta$ (Schulman, 2015), we can derive the following lower bounds:
@@
$\begin{split}\zeta(\tilde{\pi})&\geq L_{\pi}(\tilde{\pi})-\frac{2r_{\max}(1-Z)}{1-\gamma}\\ &-\frac{2\gamma\epsilon_{\mathrm{train}}}{(1-\gamma)^{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\mu_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\\ &-\frac{2\delta_{\mathrm{train}}(1-Z)}{1-\gamma}\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\\ &-\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{eval}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right],\\ \end{split}$ (9)
where $r_{\max}=\max_{s,a}\left|r(s,a)\right|$, and
$\begin{split}\delta_{\mathrm{train}}&=\max_{f\in\mathcal{F}_{\mathrm{train}}}\left\{\max_{s,a}\left|A^{\mu_{f}}(s,a)\right|\right\},\\ \delta_{\mathrm{eval}}&=\max_{f\in\mathcal{F}_{\mathrm{eval}}}\left\{\max_{s,a}\left|A^{\mu_{f}}(s,a)\right|\right\}.\\ \end{split}$ (10)
@@
See Appendix A.1. ∎
Explanation. Building on Theorems 3.2 and 3.3, we observe that, in contrast to the lower bound on training performance, the lower bound on generalization performance incorporates three additional terms, scaled by the common coefficient $(1-Z)$. This implies that increasing $Z$ contributes to improved generalization performance, with the special case of $Z=1$ resulting in alignment between generalization and training performance. Notably, this theoretical insight was also validated in Figure 2 of Cobbe et al. (2020).
@@
$\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{1}},\enspace\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{eval}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{2}},$ (11)
and
$\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\mu_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{\mathrm{train}}}.$ (12)
@@
See Appendix A.3. ∎
Therefore, $\mathfrak{D}_{1}$ can be bounded by $\mathfrak{D}_{\mathrm{train}}$. As a result, $\mathfrak{D}_{2}$ becomes crucial for improving generalization performance. Similarly, we can find an upper bound for $\mathfrak{D}_{2}$.
@@
Explanation. This definition demonstrates how the policy $\pi$ is influenced by two different rendering functions, $\tilde{f}$ and $f$, for any given underlying state $s$. If $\mathcal{R}=0$, it indicates that $D_{\mathrm{TV}}(\mu_{\tilde{f}}\|\mu_{f})[s]\equiv 0$, which means that the policy has learned a meta-representation of the observations and is no longer affected by any irrelevant features.
Figure 2: Our DML-based technique can drive agents to learn robust representations of noisy observations and gradually reduce the divergence between them, ultimately improving generalization performance.
@@
See Appendix A.5. ∎
Altogether, by combining Theorems 3.3, 3.4, 3.5, and 3.7, we can derive the following corollary.
@@
$\begin{split}&C_{\mathrm{train}}=\frac{2\delta_{\mathrm{train}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{train}}}{1-\gamma}\right)+\frac{2\gamma\epsilon_{\mathrm{train}}}{(1-\gamma)^{2}}\\ &\qquad\quad+\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right),\\ &C_{\pi}=\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right)\left(1+\frac{2\gamma\sigma_{\mathrm{train}}}{1-\gamma}\right),\\ &C_{\tilde{\pi}}=\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right),\enspace C=\frac{2r_{\max}(1-Z)}{1-\gamma}.\\ \end{split}$ (18)
@@
4 Central Hypothesis
Despite the theoretical advancements, in typical generalization settings, both the underlying Markov Decision Process (MDP) and the rendering function remain unknown. In this section, we propose that deep mutual learning (DML) (Zhang et al., 2018b) can be leveraged to enhance policy robustness against irrelevant features in high-dimensional observations, thereby improving generalization performance. This hypothesis is further illustrated in Figure 2.
The Meta-Representation Hypothesis We propose a hypothesis that deep mutual learning (DML) technique can help agents learn meta-representations of high-dimensional observations, thus improving generalization performance.

### v4

# Representation Convergence: Mutual Distillation is Secretly a Form of Regularization

@@ Abstract
In this paper, we argue that mutual distillation between reinforcement learning policies serves as an implicit regularization, preventing them from overfitting to irrelevant features. We highlight two key contributions: (a) Theoretically, for the first time, we prove that enhancing the policy robustness to irrelevant features leads to improved generalization performance. (b) Empirically, we demonstrate that mutual distillation between policies contributes to such robustness, enabling the spontaneous emergence of invariant representations over pixel inputs. Overall, our findings challenge the conventional view of distillation as merely a means of knowledge transfer, offering a novel perspective on the generalization in deep reinforcement learning. Our website: https://dml-rl.github.io/.
11footnotetext: Equal contribution. Code is available at Mutual-Distillation-Policy-Optimization.
@@ 1 Introduction
Humans exhibit a remarkable ability to learn robustly and generalize across diverse environments. Once a skill is acquired, it often transfers seamlessly to new contexts that share the same underlying semantics, even when their visual appearance differs substantially. For example, consider a person who becomes proficient at a video game, even if the background graphics or character textures are altered, the player retains their ability to perform well, effortlessly adapting to the new setting. This suggests that human learning is not overly dependent on low-level visual details, but rather grounded in abstract representations that capture the essential structure of a task. Neuroscientific studies support this view, linking abstract reasoning to the human prefrontal cortex [2, 9], and highlighting the role of inhibitory neurons in enhancing cognitive processing efficiency [31].
In stark contrast, visual reinforcement learning (VRL) agents often struggle with generalization. While they can be trained to solve complex tasks in specific environments, even minor changes, such as shifts in color schemes or background textures, can significantly degrade their performance. This sensitivity indicates that VRL agents tend to overfit to superficial visual features, failing to capture the underlying structure of the task [4, 5]. These limitations give rise to a fundamental question:
What hinders reinforcement learning agents from generalizing like humans? How can we enable them to learn robust representations that drive human-like generalization behavior?
The core reason behind the limited generalization ability of VRL agents lies in their reliance on convolutional neural networks (CNNs) as visual encoders. While CNNs are the de facto choice for processing high-dimensional visual inputs, they are notoriously sensitive to even small perturbations [13]. This brittleness significantly hampers the robustness of learned policies and limits their ability to generalize across visually diverse environments. To address this issue, one common strategy is to apply data augmentation [39], which improves robustness by diversifying the training distribution and reducing dataset-induced biases. Alternatively, invariant representation learning has emerged as a principled approach to tackle generalization problem from a feature-learning perspective. It aims to extract representations that remain stable under a wide range of input transformations, thereby promoting robustness and transferability [29].
While data augmentation is an effective bias mitigation technique, its reliance on task-specific strategies that are manually crafted by human experts, poses a challenge for designing task-independent solutions. In contrast, our method enables agents to generalize without any handcrafted augmentations or external priors, relying purely on training experience. Invariant representation learning is a promising approach to enhance model’s cross-domain generalization. However, it relies on transformation correspondences, which are fundamentally inaccessible in the generalization scenarios of reinforcement learning due to the dynamic nature of environments. In addition, invariant representation framework inherently separates the encoder from the model, unnecessarily complicating the theoretical analysis. Instead, our framework is theoretically and empirically end-to-end.
In this paper, we first propose a novel theoretical framework to analyze the generalization problem in reinforcement learning and show that the policy robustness to irrelevant features enhances its generalization performance. Building upon this principled insight, we then provide empirical evidence that deep mutual learning (DML) [53] can implicitly prevent them from overfitting to such irrelevant features, leading to robust learning process and significant generalization improvements.
In summary, the main contributions of this paper are as follows:
@@
  • We theoretically prove that improving the policy robustness to irrelevant features enhances its generalization performance. To the best of our knowledge, we are the first to provide a rigorous proof of this intuition.
@@
  • We propose a hypothesis that deep mutual learning (DML) enhances the generalization performance of the policy by implicitly regularizing irrelevant features. We also provide intuitive insights to support this hypothesis.
@@ 2.1 Markov Decision Process and Generalization
$Q^{\mu}(s_{t},a_{t})=\mathbb{E}_{\mu}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k})\right],\enspace V^{\mu}(s_{t})=\mathbb{E}_{a_{t}\sim\mu(\cdot|s_{t})}\left[Q^{\mu}(s_{t},a_{t})\right].$ (1)
@@
Explanation. Assumption 2.2 states that all $\mathcal{M}_{f}$ share a common underlying MDP $\mathcal{M}$.
@@
where $o_{0}^{f}\sim\rho_{f}(\cdot)$, $a_{t}\sim\pi(\cdot|o_{t}^{f})$, $r_{t}^{f}=r_{f}(o_{t}^{f},a_{t})$ and $o_{t+1}^{f}\sim\mathcal{P}_{f}(\cdot|o_{t}^{f},a_{t})$, we simplify the notation to $\tau_{f}\sim\pi$. During training, the agent is only allowed to access a subset of all MDPs, which is $\left\{\mathcal{M}_{f}|f\in\mathcal{F}_{\mathrm{train}}\subset\mathcal{F}\right\}$, and then tests its generalization performance across all MDPs. Thus, denote $p_{\mathrm{train}}:\mathcal{F}_{\mathrm{train}}\mapsto[0,1]$ as the distribution over $\mathcal{F}_{\mathrm{train}}$, the agent’s training performance $\eta(\pi)$ and generalization performance $\zeta(\pi)$ can be expressed as
$\eta(\pi)=\mathbb{E}_{f\sim p_{\mathrm{train}}(\cdot),\tau_{f}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{f}(o_{t}^{f},a_{t})\right],\enspace\zeta(\pi)=\mathbb{E}_{f\sim p(\cdot),\tau_{f}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{f}(o_{t}^{f},a_{t})\right].$ (4)
@@ 2.2 Deep Mutual Learning
Deep mutual learning (DML) [53] is a mutual distillation technique in supervised learning. Unlike the traditional teacher-student distillation strategy, DML aligns the probability distributions of multiple student networks by minimizing the KL divergence loss during training, allowing them to learn from each other. Specifically,
@@
where $\mathcal{L}_{\mathrm{SL}}$ and $\mathcal{L}_{\mathrm{KL}}$ represent the supervised learning loss and the KL divergence loss, respectively, $\alpha$ is the weight. Using DML, the student cohort effectively pools their collective estimate of the next most likely classes. Finding out and matching the other most likely classes for each training instance according to their peers increases each student’s posterior entropy, which helps them converge to a more robust representation, leading to better generalization.
@@ 3 Theoretical Results
$p_{\mathrm{train}}(f)=\frac{p(f)\cdot\mathbb{I}(f\in\mathcal{F}_{\mathrm{train}})}{Z},\enspace p_{\mathrm{eval}}(f)=\frac{p(f)\cdot\mathbb{I}(f\in\mathcal{F}_{\mathrm{eval}})}{1-Z},$ (6)
where $Z=\int_{\mathcal{F}_{\mathrm{train}}}p(f)\mathrm{d}f$ and $1-Z$ is the normalization term, $\mathcal{F}_{\mathrm{eval}}=\mathcal{F}-\mathcal{F}_{\mathrm{train}}$, $\mathbb{I}(\cdot)$ denotes the indicator function.
@@
We can thus analyze the generalization problem using the underlying policy $\mu_{f}$. Then, define $L_{\pi}$ as the first-order approximation of $\eta$ [36], we can derive the following lower bounds:
@@
where $r_{\max}=\max_{s,a}\left|r(s,a)\right|$, $\delta_{\mathrm{train}}=\max_{f\in\mathcal{F}_{\mathrm{train}}}\left\{\max_{s,a}\left|A^{\mu_{f}}(s,a)\right|\right\}$, and $\delta_{\mathrm{eval}}=\max_{f\in\mathcal{F}_{\mathrm{eval}}}\left\{\max_{s,a}\left|A^{\mu_{f}}(s,a)\right|\right\}$.
@@
See Appendix G.1. ∎
Explanation. Building upon Theorems 3.2 and 3.3, we observe that, in contrast to the lower bound on training performance, the lower bound on generalization performance incorporates three additional terms, scaled by the common coefficient $(1-Z)$. This implies that increasing $Z$ contributes to improved generalization performance, with the special case of $Z=1$ resulting in alignment between generalization and training performance. Notably, this theoretical insight was also validated in Figure 2 of Cobbe et al. [5].
@@
$\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{1}},\enspace\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{eval}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{2}},\enspace\underbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\mu_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]}_{\text{denote it as }\mathfrak{D}_{\mathrm{train}}}.$ (10)
@@
See Appendix G.3. ∎
Therefore, $\mathfrak{D}_{1}$ can be bounded by $\mathfrak{D}_{\mathrm{train}}$. As a result, $\mathfrak{D}_{2}$ becomes crucial for improving generalization performance. Similarly, we can find an upper bound for $\mathfrak{D}_{2}$:
@@
Explanation. This definition demonstrates how the policy $\pi$ is influenced by two different rendering functions, $\tilde{f}$ and $f$, for any given underlying state $s$. If $\mathcal{R}=0$, it indicates that $D_{\mathrm{TV}}(\mu_{\tilde{f}}\|\mu_{f})[s]\equiv 0$, which means that the policy is no longer affected by any irrelevant features.
@@
See Appendix G.5. ∎
Altogether, by combining Theorems 3.3, 3.4, 3.5, and 3.7, we can derive the following corollary:
@@
$\begin{split}C_{\mathrm{train}}&=\frac{2\delta_{\mathrm{train}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{train}}}{1-\gamma}\right)+\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right)+\frac{2\gamma\epsilon_{\mathrm{train}}}{(1-\gamma)^{2}},\\ C_{\pi}&=\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right)\left(1+\frac{2\gamma\sigma_{\mathrm{train}}}{1-\gamma}\right),\\ C_{\tilde{\pi}}&=\frac{2\delta_{\mathrm{eval}}(1-Z)}{1-\gamma}\left(1+\frac{2\gamma\sigma_{\mathrm{eval}}}{1-\gamma}\right),\enspace C=\frac{2r_{\max}(1-Z)}{1-\gamma}.\\ \end{split}$ (16)
Figure 1: This is a toy environment where the gray agent’s goal is to pick up coins.
@@
4 Mutual Distillation is Secretly a Form of Regularization
Despite the theoretical advancements, in typical generalization settings, both the underlying MDP and the rendering function remain unknown. Next, we begin by introducing a minimal toy example in Section 4.1, which we then provide an in-depth analyze in Section 4.2 to motivate our hypothesis.
4.1 Toy Example
Let’s consider a simple environment where the agent attempts to pick up coins to earn rewards (see Figure 1). The agent’s observations are the current pixels. It is clear that the agent’s true objective is to pick up the coins, and the background color is a spurious feature. However, upon observing the training data for policy A, we can see that in the red background, the coins are always on the right side of the agent, while in the cyan background, the coins are always on the left side. As a result, when training policy A using reinforcement learning algorithms, it is likely to exhibit overfitting behavior, such as moving to the right in a red background and to the left in a cyan background.
However, the overfitting of policy A to the background color will fail in the training data of policy B, because in policy B’s training data, regardless of whether the background color is red or cyan, the coin can appear either on the left or right side of the agent. Therefore, through DML, policy A is regularized by the behavior of policy B, effectively preventing policy A from overfitting to the background color. In other words, any irrelevant features learned by policy A could lead to suboptimal performance of policy B, and vice versa. Thus, we hypothesize that this process will force both policies to learn the true underlying semantics, ultimately improving generalization performance.
Figure 2: (Left) Independently trained reinforcement learning policies may overfit to spurious features. (Right) Through mutual distillation via DML, two policies regularize each other to converge toward a more robust hypothesis space, ultimately improving generalization performance.
4.2 Central Hypothesis